### v1

@@ Appendix C Pre-processing results
Appendix D Additional results on gastroenterology exam
Additional results related to the flexible endoscope colonoscopy with possible biopsies examination are reported in this section. The dataset for this exam consisted of all referrals filled in in the Lombardy Region between 2019 and 2021, for a total of 407,949 referrals. A random sample of 1,000 distinct clinical questions (corresponding to 3,556 referrals) was annotated in the same way as for the case study presented in the main manuscript and used to validate the results.
Considering 2015 guidelines from Regione Lombardia for this type of exam [27], we marked colon cancer familiarity/prevention and chronic inflammatory bowel disease as likely appropriate. Guidelines state that, in these cases, physicians should carefully consider the patient history and condition to establish if and when the exam should be performed.
Table D.1 reports the detailed results per category, while Table D.2 the summary metrics considering the categories and the appropriate/inappropriate levels. Table D.3 presents the final results on the full dataset, which are stratified by year, type of prescribing physician and LHU in Figure D.1.
Table D.4 details the labels from manual annotation on the test set.
Table D.1: Detailed results of the cluster mapping for each category, on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score.
Category P-CQ [%] R-CQ [%] F1-CQ [%] # CQ P-REF [%] R-REF [%] F1-REF[%] # REF
Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
Table D.2: Summary metrics on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF), considering annotations categories or only appropriate/inappropriate/likely appropriate/undetermined. P=Precision, R=Recall, F1=F1-Score.
P-CQ [%] R-CQ [%] F1-CQ [%] P-REF [%] R-REF [%] F1-REF [%]
Annotation categories 90.42 88.78 88.60 93.59 92.70 92.67
Appropriateness levels 92.05 90.96 90.88 94.66 93.96 94.05
Table D.3: Summary of referrals’ clusters for flexible endoscope colonscopy, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
Cluster Label # REF %
polyps 81,111 19.88
occult blood 76,487 18.75
colon cancer familiarity / prevention 43,615 10.69
abdominalgia 31,316 7.68
colon cancer 30,434 7.46
anaemia 23,133 5.67
@@ Appendix D Additional results on gastroenterology exam
chronic inflammatory bowel disease 21,470 5.26
constipation 15,109 3.70
diverticulosis 15,058 3.69
chronic diarrhoea 13,533 3.32
altered alvo 10,933 2.68
haemorrhoidal disease 5,379 1.32
diverticulitis 4,669 1.14
suspected colic lesion detected by other investigation technique 3,671 0.90
irritable bowel syndrome 2,023 0.50
surgery 797 0.20
lynch syndrome 787 0.19
other 28,424 6.98
Figure D.1: Prescription appropriateness for flexible endoscope colonscopy grouped by LHA (A), type of prescribing physicians (B) and year (C)
Table D.4: Labels derived by manual annotations on the validation set of 1000 clinical questions for flexible endoscope colonoscopy.
Label # CQ # REF
polyps 158 344
bleeding 122 954
colon cancer 107 187
abdominalgia 104 214
colon cancer familiarity 82 141
anaemia 65 115
prevention 57 161
occult blood 36 789
constipation 36 139
diverticulosis 33 48
altered bowel 30 65
colitis 28 40
chronic diarrhoea 28 107
diverticulitis 14 28
crohn’s disease 12 16
haemorrhoidal disease 9 30

### v2

@@ Appendix B Pre-processing results
Table B.3: Results of clusters mapped to the disease labels on the manually labelled test sets, computed for clinical question (CQ) and referrals (REF) without applying pre-processing steps. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
ECD
Umberto-E3C-Pipeline 84.01 73.51 78.40 89.15 79.28 83.95
Umberto-Base 77.99 70.68 74.18 85.01 73.18 78.66
Word2Vec 75.63 62.25 68.27 78.90 67.77 72.99
TF-IDF 46.28 33.61 38.92 45.73 37.83 41.46
LDA 24.38 32.27 27.88 16.51 25.97 20.24
FEC
Umberto-E3C-Pipeline 88.59 85.53 87.12 90.99 87.65 89.32
Umberto-Base 80.28 70.31 75.02 85.79 75.42 80.35
Word2Vec 78.66 68.55 73.27 80.15 70.38 74.96
TF-IDF 55.45 44.91 49.69 55.95 49.60 52.65
LDA 31.52 36.30 33.80 26.70 33.91 29.86
Appendix C Clustering
C.1 Hyperparameters and algorithms selection
Table LABEL:tab:hdbscan_hyper reports the grid search for HDBSCAN hyperparameters, evaluated with F1-CQ on ECD dataset.
Table C.1: Grid search for HDBSCAN hyperparameters, with F1-CQ on ECD dataset
min_cluster_size min_samples cluster_selection_epsilon F1-CQ [%]
20 0 11.42
20 0.1 11.42
20 0.5 11.42
20 1 0 63.91
20 1 0.1 63.91
20 1 0.5 63.91
20 5 0 10.13
20 5 0.1 10.13
20 5 0.5 10.13
20 10 0 77.85
20 10 0.1 77.85
20 10 0.5 77.85
20 20 0 11.42
20 20 0.1 11.42
20 20 0.5 11.42
50 0 40.31
50 0.1 40.31
50 0.5 40.31
50 1 0 60.81
50 1 0.1 60.81
50 1 0.5 60.81
50 5 0 71.72
50 5 0.1 71.72
50 5 0.5 71.72
50 10 0 70.40
50 10 0.1 70.40
50 10 0.5 70.40
50 20 0 40.31
50 20 0.1 40.31
50 20 0.5 40.31
100 0 23.87
100 0.1 23.87
100 0.5 23.87
100 1 0 75.64
100 1 0.1 75.64
100 1 0.5 75.64
100 5 0 44.60
100 5 0.1 44.60
100 5 0.5 44.60
100 10 0 51.82
100 10 0.1 51.82
100 10 0.5 51.82
100 20 0 40.31
100 20 0.1 40.31
100 20 0.5 40.31
200 0 14.46
200 0.1 14.46
200 0.5 14.46
200 1 0 67.41
200 1 0.1 67.41
200 1 0.5 67.41
200 5 0 36.43
200 5 0.1 36.43
200 5 0.5 36.43
200 10 0 43.32
200 10 0.1 43.32
200 10 0.5 43.32
200 20 0 31.80
200 20 0.1 31.80
200 20 0.5 31.80
500 0 18.64
500 0.1 18.64
500 0.5 18.64
500 1 0 15.48
500 1 0.1 15.48
500 1 0.5 15.48
500 5 0 14.45
500 5 0.1 14.45
500 5 0.5 14.45
500 10 0 23.37
500 10 0.1 23.37
500 10 0.5 23.37
500 20 0 16.31
500 20 0.1 16.31
500 20 0.5 16.31
step from 500 0 34.52
step from 500 0.1 34.52
step from 500 0.5 34.52
step from 500 1 0 71.54
step from 500 1 0.1 71.54
step from 500 1 0.5 71.54
step from 500 5 0 79.48
step from 500 5 0.1 79.48
step from 500 5 0.5 79.48
step from 500 10 0 63.28
step from 500 10 0.1 63.28
step from 500 10 0.5 63.28
step from 500 20 0 41.95
step from 500 20 0.1 41.95
step from 500 20 0.5 41.95
Table C.1: *
Figures C.1 and C.2 report the performance of K-Means and Agglomerative Hierarchical Clustering on the ECD dataset, at the CQ level, varying the number of clusters. Their results are consistently lower than those from HDBSCAN.
Figure C.1: K-means performance on ECD-CQ varying the number of clusters
Figure C.2: Agglomerative hierarchical clustering performance on ECD-CQ varying the number of clusters
For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens. The best results presented are obtained with K-Means. The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40 for ECD and k=35 for FEC.
For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7 for ECD and k=9 for FEC.
For direct embeddings match, we matched Umberto embeddings of clinical questions with embeddings of known diseases/symptoms from guidelines, using cosine similarity. If no element in the guidelines matched with a similarity of at least 0.6 we assigned the Other label.
C.2 Additional clustering results
Tables LABEL:tab:clusters-summ and LABEL:tab:clusters-summ-fec present all the clusters identified for ECD and FEC, respectively, with their keyword descriptions, the sizes and the labels they were assigned.
Table C.2: Summary of all clusters identified for ECD, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
1 claudicatio; riferita claudication; referred claudication 1132 1686
2 claudicatio; intermittens claudicatio; intermittens claudication 653 1179
3 claudicatio claudicatio claudication 402 530
4 claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudication 60 126
5 claudicatio; arti; inferiori claudicatio; limbs; lower claudication 41 49
6 dermatite dermatitis dermatitis 421 474
7 diabete diabetes diabetes 1540 3249
8 diabete; mellito diabetes; mellitus diabetes 534 1847
9 diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
10 diabetico; diabetica diabetes; diabetic diabetes 209 165
11 dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
12 dialisi dialysis dialysis 110 222
13 dimero dimer dimer 1343 1601
14 dispnea dyspnoea dyspnoea 104 171
15 dolore pain pain 6343 12051
16 deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
17 algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
18 polpaccio; dolore calf; pain pain 707 1753
19 algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
20 dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
21 dolore; popliteo pain; popliteal pain 35 117
22 algia; covid; gamba pain; covid; leg pain 30 63
23 dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
24 crampi; notturni cramps; nocturnal pain 6 5
25 edema edema edema 11686 20723
26 edemi; declivi edema; declivities edema 9491 16473
27 edema; arti; edemi edema; limbs; edemas edema 3712 7867
28 arti; inferiori; edemi limbs; lower; edemas edema 528 2486
29 edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
30 edema; arti; inferiori edema; limbs; lower edema 899 2350
31 edema; perimalleolare edema; perimalleolar edema 1039 2214
32 edema; natura; determinare edema; nature; determine edema 1027 2199
33 edemi; perimalleolari edema; perimalleolar edema 1190 2151
34 edema; malleolare edema; malleolar edema 935 1136
35 edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
36 edema; flebite edema; phlebitis edema 217 297
37 arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
38 arti; inferiori; edema limbs; lower; edema edema 80 168
39 arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
40 edema; inferiori; arti edema; lower; limbs edema 74 163
41 edema; sinistra edema; left edema 61 108
42 claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
43 linfedema; arti lymphedema; limbs edema 23 24
44 emorragia haemorrhage haemorrhage 18 29
45 formicolio tingling tingling 43 170
46 freddo cold cold 49 147
47 fumatore smoker smoker 121 193
48 gravidanza pregnancy pregnancy 92 225
49 insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
50 venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
51 venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
52 superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
53 insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
54 venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
55 venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
56 venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
57 insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
58 venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
59 venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
60 insufficienza; venosa venous; insufficiency venous insufficiency 0 0
61 arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
62 venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
63 venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
64 venosa; cronica venous; chronic venous insufficiency 0 0
65 insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
66 intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
67 lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
68 lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
69 ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
70 ipoestesia hypoesthesia hypoesthesia 31 80
71 ischemia ischaemia ischaemia 517 617
72 parestesie paresthesias paresthesias 1834 3965
73 parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
74 pesantezza heaviness heaviness 215 475
75 arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
76 arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
77 safenectomia saphenectomy saphenectomy 1988 4177
78 safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
79 safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
80 safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
81 sindrome metabolica metabolic syndrome metabolic syndrome 12 48
82 stent stent stent 301 612
83 tromboflebite thrombophlebitis thrombophlebitis 7073 17415
84 flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
85 flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
86 varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
87 tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
88 gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
89 tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
90 flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
91 flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
92 tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
93 flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
94 tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
95 trombosi thrombosis thrombosis 17060 43558
96 trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
97 trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
98 trombosi; venosa venous; thrombosis thrombosis 1325 3337
99 venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
100 poplitea; trombosi popliteal; thrombosis thrombosis 607 642
101 venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
102 trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
103 segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
104 safena; trombosi saphenous vein; thrombosis thrombosis 301 263
105 post; trombotica post; thrombotic thrombosis 36 73
106 tumefazione swelling swelling 533 741
107 tumefazione; gamba swelling; leg swelling 172 448
108 destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
109 ulcera; vascolare ulcer; vascular ulcer 828 3861
110 ulcera ulcer ulcer 1258 2411
111 varici varices varices 13999 35939
112 arti; inferiori; varici limbs; lower; varices varices 7018 18157
113 venose; varici venous; varices varices 4957 13554
114 varicosita; stasi varicosities; stasis varices 957 2701
115 varici; arti; inferiori varices; limbs; lower limbs varices 219 277
116 vene; varicose; algie veins; varicose; algie varices 121 194
117 safena; varici saphenous veins; varicose varices 47 136
118 varicosita; varicosita’ varicosities; varicositis varices 86 97
119 vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
120 varici; sospette varicose veins; suspected varices 34 63
121 varicosita; gamba varicosities; leg varices 80 63
122 varici; intervento varicosities; surgery varices intervention 3144 7126
123 vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
124 vasculopatia vasculopathy; peripheral vasculopathy 769 1886
125 vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
126 arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
127 arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
128 iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
129 aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
130 vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
131 vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
132 vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
133 arteriopatia arteriopathy vasculopathy 20 25
134 natura; determinare; periferica nature; determine; peripheral other 3978 10039
135 natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
136 paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
137 arti; inferiori limbs; lower other 2190 5165
138 safena; incontinenza saphenous; incontinence other 288 782
139 ipertrofia hypertrophy hypertrophy 506 591
140 tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
141 embolia; polmonare embolism; pulmonary embolism 110 174
142 femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
143 stenosi; venoso stenosis; venous other 30 49
144 dislipidemia; ipertensione dyslipidaemia; hypertension dyslipidaemia 25 44
145 aorta; sottorenale aorta; subrenal other 25 29
Table C.2: *
Table C.3: Summary of all clusters identified for FEC, including their keyword representation, the label assigned to them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
1 addominalgia; alvo abdominalgia; alvo abdominal pain 163 371
2 addominalgie; natura; determinare abdominalgia; nature; determine abdominal pain 714 5.801
3 addominali; algie; dolori abdominalgia; algia; pain abdominal pain 894 5.430
4 addominali; coliche abdominal; colic abdominal pain 240 3.775
5 alvo; addominalgie bowel; abdominal pain abdominal pain 224 2.507
6 calo; ponderale; addominalgia fall; weight; abdominalgia abdominal pain 100 215
7 natura; determinare; coliche nature; determine; colic abdominal pain 325 2.590
8 natura; determinare; addominali; coliche nature; determine; abdominal; colic abdominal pain 3.687 5.605
9 fianco; dolore; persistente flank; pain; persistent abdominal pain 61 121
10 fossa; addominalgia fossa; abdominalgia abdominal pain 192 2.125
11 fossa; iliaca; algia fossa; iliac; algia abdominal pain 90 194
12 fossa; iliaca; algie fossa; iliac; algia abdominal pain 33 71
13 fossa; iliaca; dolore fossa; iliac; pain abdominal pain 302 662
14 ipocondrio; destra; dolore; sinistra hypochondrium; right; pain; left abdominal pain 127 281
15 quadranti; addominalgia; addominalgie quadrant; abdominalgia; abdominalgia abdominal pain 33 74
16 stipsi; addominalgia constipation; abdominalgia abdominal pain 501 1.105
17 stipsi; addominali constipation; abdominal abdominal pain 112 250
18 stipsi; addominali; algie constipation; abdominal; pain abdominal pain 61 139
19 alvo; alterazione; alterazioni alvo; alteration altered bowel 432 3.500
20 alvo; alterno bowel; abdominalgia altered bowel 610 2.496
21 alvo; alterno; addominalgia bowel; abdominalgia altered bowel 287 1.925
22 alvo; calo; ponderale bowel; decline; weight altered bowel 196 1.655
23 alvo; stipsi bowel; constipation altered bowel 129 277
24 fossa; alvo fossa; alvo altered bowel 197 1.080
25 anemia; emoglobina anaemia; haemoglobin anaemia 61 128
26 anemia; natura; determinare; grave anaemia; nature; determine; severe anaemia 602 4.392
27 anemia; sangue; occulto; positivo anaemia; blood; occult; positive anaemia 315 2.366
28 anemia; sideropenica; origine; indeterminata anaemia; sideropenic; origin; undetermined anaemia 332 6.663
29 anemia; sideropenica; sangue; occulto anaemia; sideropenic; blood; occult anaemia 49 112
30 anemizzazione; grave; severa; sangue anaemia; severe; blood anaemia 456 3.814
31 anemizzazione; natura; determinare anaemia; nature; determine anaemia 250 2.504
32 sideropenica; anemia iron deficiency; anaemia anaemia 365 3.154
33 diarrea; colite; cronica diarrhoea; colitis; chronic chronic diarrhoea 95 757
34 diarrea; cronica diarrhoea; chronic chronic diarrhoea 778 7.668
35 diarroico; alvo diarrhoeic; alvo chronic diarrhoea 193 2.694
36 natura; determinare; diarrea; dolore; addominale nature; determine; diarrhoea; pain; abdominal chronic diarrhoea 409 2.414
@@ C.2 Additional clustering results
37 colite; ischemica colitis; ischaemic chronic inflammatory bowel disease 119 286
38 crohn; morbo crohn disease; disease chronic inflammatory bowel disease 540 3.812